### v1

# Rep3D: Re-parameterize Large 3D Kernels with Low-Rank Receptive Modeling for Medical Imaging

Abstract
In contrast to vision transformers, which model long-range dependencies through global self-attention, large kernel convolutions provide a more efficient and scalable alternative, particularly in high-resolution 3D volumetric settings. However, naïvely increasing kernel size often leads to optimization instability and degradation in performance. Motivated by the spatial bias observed in effective receptive fields (ERFs), we hypothesize that different kernel elements converge at variable rates during training. To support this, we derive a theoretical connection between element-wise gradients and first-order optimization, showing that structurally re-parameterized convolution blocks inherently induce spatially varying learning rates. Building on this insight, we introduce Rep3D, a 3D convolutional framework that incorporates a learnable spatial prior into large kernel training. A lightweight two-stage modulation network generates a receptive-biased scaling mask, adaptively re-weighting kernel updates and enabling local-to-global convergence behavior. Rep3D adopts a plain encoder design with large depthwise convolutions, avoiding the architectural complexity of multi-branch compositions. We evaluate Rep3D on five challenging 3D segmentation benchmarks and demonstrate consistent improvements over state-of-the-art baselines, including transformer-based and fixed-prior re-parameterization methods. By unifying spatial inductive bias with optimization-aware learning, Rep3D offers an interpretable, and scalable solution for 3D medical image analysis. The source code is publicly available at https://github.com/leeh43/Rep3D.
1 Introduction
The landscape of medical vision models has evolved rapidly, expanding from early convolutional architectures to modern transformer-based designs. In particular, Vision Transformers (ViTs) have gained traction for their ability to model long-range dependencies using multi-head self-attention and minimal inductive bias [9]. In parallel, the community has revisited large kernel convolutions as a scalable alternative to attention mechanisms, particularly in the context of high-resolution 3D volumetric data [24, 18]. Despite architectural differences, both ViTs and large-kernel CNNs share a central goal: expanding the effective receptive field (ERF) to enable rich spatial context aggregation. However, simply increasing kernel size does not guarantee improved performance. Prior work has shown that naïve enlargement of convolutional filters can result in saturated or degraded accuracy across various segmentation tasks [7, 19]. Unlike ViTs, which adaptively attend to spatial content, standard convolutions rely on static, weight-shared kernels and lack the ability to modulate importance across spatial positions. This limitation prompts our first research question: Can we incorporate spatial priors into large kernel convolutions to improve learning effectiveness?
Recent advances in structural re-parameterization offer a promising direction. Methods such as RepLKNet [7], SLaK [22], and PELK [3] scale kernels to extreme sizes (e.g., $31\times 31$, $51\times 51$, $101\times 101$) by combining parallel branches of “large + small” convolutions into what is referred to as a Constant-Scale Linear Addition (CSLA) block. These parallel paths are merged into a single kernel at inference time, enabling efficient deployment while capturing multi-scale features during training. Interestingly, we observe that CSLA blocks naturally encode spatial learning bias: elements near the kernel center tend to converge faster than those on the periphery. This mirrors diffusion-like gradient propagation in ERFs starting from the center and expanding outward. These observations suggest that convergence dynamics are not uniform across the kernel, but instead spatially structured. This leads to our second question: Can we explicitly model this diffusion pattern as a learnable spatial prior to re-weight kernel element updates during training?
To address this, we first provide a theoretical analysis of the optimization dynamics in CSLA-based re-parameterized convolutions. We show that each branch (e.g., small vs. large kernels) can implicitly operates under a distinct learning rate, leading to element-wise differences in convergence speed. These dynamics correlate with ERF visualizations and share characteristics with spatial frequency patterns in human visual perception [17]. Inspired by this, we propose a novel receptive bias re-parameterization strategy that encodes spatial distance from the kernel center as a spatial bias prior on learning convergence. We implement this as a low-rank modulation mechanism that generates spatial scaling factors for kernel weights, allowing the optimizer to emphasize local versus global regions adaptively for gradient back-propagation.
Building on this insight, we present Rep3D, a 3D convolutional architecture that integrates large kernel convolutions (e.g., $21\times 21\times 21$) with our proposed re-parameterization approach. Unlike prior approaches that rely on multi-branch structures, Rep3D employs a plain and efficient encoder to reduce complexity while preserving representational capacity. We evaluate Rep3D across five challenging volumetric medical segmentation benchmarks and show that it consistently outperforms state-of-the-art transformer- and CNN-based models.Our key contributions are as follows:
  • We propose Rep3D, a 3D CNN with large kernel convolutions and a streamlined encoder design that achieves state-of-the-art (SOTA) performance on multi-scale (i.e. from organs/tissues to tumors) segmentation benchmarks.
  • We propose a novel and theoretically grounded re-parameterization approach that models ERF diffusion as a learnable spatial bias prior, enabling element-wise modulation of gradient convergence for training.
  • We validate our method on five challenging 3D medical imaging benchmarks under direct training settings, achieving consistent and significant improvements across all datasets.
Figure 1: (a) Traditional structural re-parameterization methods (e.g., CSLA blocks) re-parameterize small and large kernel convolutions to improve representational capacity, but apply linear optimization with same learning rate across the kernels, demonstrating a faster convergence in local then global. (b) In contrast, Rep3D introduces a learnable spatial bias via a generator network $f_{\theta}$, which modulates each element in the large kernel using a prior based on distance decay. This adaptive modulation enables local-to-global update dynamics aligned with ERF behavior, enhancing both training stability and model performance for 3D volumetric tasks.
2 Related Work
The Transition from CNN to ViT. Convolutional neural networks (CNNs) have long served as the foundation for medical image segmentation, with the U-Net architecture [27, 5] establishing a dominant encoder–decoder paradigm for dense prediction. Variants such as V-Net [25], UNet++ [33], H-DenseUNet [21], and SegResNet [26] extended this architecture to better suit 2D and 3D modalities, as well as different anatomical contexts. More recently, nnU-Net [14] automated the design of 3D segmentation pipelines with a coarse-to-fine strategy tailored for various medical datasets. However, most CNN-based architectures rely on small kernel sizes, limiting their effective receptive field (ERF) and making it difficult to capture long-range dependencies. Vision Transformers (ViTs), starting with TransUNet [4], brought global attention mechanisms to medical segmentation, allowing models to attend across distant voxels. Models such as UNETR [11], CoTr [29], and LeViT-UNet [30] have shown strong performance, particularly in organ and tumor segmentation. Yet, the quadratic complexity of self-attention poses major bottlenecks for high-resolution volumetric data. To mitigate this, hierarchical transformers such as Swin Transformer [23] introduced localized attention via a sliding window mechanism. Follow-up models like SwinUNETR [10], nnFormer [32], Swin-Unet [2], and SwinBTS [16] adapted this approach to efficiently model multi-scale features. Despite these innovations, transformer-based methods remain computationally expensive and slow to train, particularly for dense 3D segmentation tasks. Meanwhile, depthwise large-kernel CNNs (e.g., ConvNeXt [24]) offer a promising compromise by mimicking ViT-like receptive fields with fewer computational demands. 3D UX-Net [18] and similar designs apply this principle to volumetric medical data, although challenges remain in segmenting fine-scale anatomical structures or multi-scale lesions without additional spatial priors or architectural adaptations.
The integration of Weight Re-parameterization. Structural re-parameterization (SR) has emerged as a powerful paradigm to enhance CNN training without altering inference-time complexity. Models like RepVGG [8] and OREPA [13] employ additional convolution branches (e.g., $1\times 1$ or identity paths) during training to improve gradient flow and feature diversity. These branches are merged into a single convolution kernel post-training, allowing for efficient inference. RepLKNet [7] and SLaK [22] extend this approach to large 2D kernels (e.g., $31\times 31$ and $51\times 51$), increasing the receptive field while maintaining tractable inference cost via kernel decomposition or sparse groups. A complementary line of work focuses on gradient re-parameterization instead of modifying model weights directly. RepOptimizer [6], for example, modifies the back-propagation process by applying learnable scaling to gradient updates, enabling effective training of plain CNNs. These techniques reduce reliance on complex architectural design and have been shown to match or exceed the performance of more intricate networks. While much of the re-parameterization research has focused on 2D natural images, extending these methods to 3D medical imaging presents unique challenges. Volumetric kernels require significantly more parameters, and naive kernel expansion leads to high computational costs and optimization instability. 3D RepUX-Net [19] demonstrate the inital attempt of adapting weight re-parameterization to 3D medical imaging and scale large depthwise kernels with fixed prior context, but still lacks of flexibility on adapting dynamic variation in fine-grained semantics for learning convergence. To bridge this gap, there is growing interest in using spatial priors or effective receptive field modeling to guide re-parameterization for large kernel learning in the 3D setting.
3 Rep3D
Figure 2: In contrast to (a) structural or (b) gradient-based re-parameterization, Rep3D introduces a novel re-parameterization strategy that injects a learnable spatial bias into large kernel convolutions for optimization. During training, a lightweight generator network produces a modulation mask conditioned on a distance-based prior, which adaptively scales gradient updates across the kernel. This enables spatially-aware learning dynamics that reflect local-to-global variations in the effective receptive field (ERF).
Rep3D rethinks the training dynamics of large-kernel convolution by explicitly embedding spatial bias, derived from effective receptive fields (ERFs), into the optimization process. Motivated by structural reparameterization (SR) and the distinctive gradient behavior observed in ERFs, Rep3D introduces a low-rank, learnable reparameterization that adapts element-wise update behavior across the kernel. We first derive the theoretical equivalence between parallel convolution branches and their single-operator counterparts, showing that a “large + small” convolution block (as in RepLKNet [7]) implicitly assigns spatially varying learning rates. We then translate this insight into a unified formulation and construct a lightweight generator that outputs a convergence-aware modulation mask. The output modulated mask models fine-grained learning dynamics during training, improving both scalability and performance in 3D tasks with large kernel convolution.
3.1 Variable Learning Convergence in Parallel Branch
As shown in Figure 2, the learning convergence of the large kernel convolution can be improved by either adding up the encoded outputs of parallel branches weighted by diverse scales with SR (RepLKNet [7]) or performing Gradient Reparameterization (GR) by multiplying with constant values (RepOptimizer [6]) in a Single Operator (SO). Inspired by the concepts of SR and GR, we extend the theoretical derivation from RepOptimizer and observe the variable learning rate across branches. We begin by analyzing the CSLA block, a basic two-branch design used in SR-based networks (e.g., RepLKNet [7]). Let $X$ denote the input feature map, and let $W_{L},W_{S}$ be large and small 3D convolution kernels, scaled by fixed positive scalars $\alpha_{L}$ and $\alpha_{S}$, respectively. The output of the CSLA module is:
$Y_{\text{CSLA}}=\alpha_{L}(X\ast W_{L})+\alpha_{S}(X\ast W_{S}),$ (1)
where $\ast$ denotes 3D convolution. To unify the branches into a single equivalent convolution for efficient inference, we define a single-operator (SO) form:
$Y_{\text{SO}}=X\ast W^{\prime},$ (2)
where the equivalent kernel $W^{\prime}$ is a linear combination of the two branches:
$W^{\prime}=\alpha_{L}W_{L}+\alpha_{S}W_{S}.$ (3)
During training with first-order optimization (i.e. SGD, AdamW) and step size $\lambda$, we apply the stochastic gradient descent rule and update the gradients for the parallel branches as follow:
$W^{\prime}_{t+1}=W^{\prime}_{t}-\lambda\frac{\partial\mathcal{L}}{\partial W^{\prime}_{t}}.$ (4)
As the parallel branch architecture updates $W_{L}$ and $W_{S}$ independently:
$W_{L(t+1)}$ $=W_{L(t)}-\lambda_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}},\quad W_{S(t+1)}=W_{S(t)}-\lambda_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}.$ (5)
where $\lambda_{L}$ and $\lambda_{S}$ are the learning rate for corresponding branch respectively. Substituting these into the equivalent kernel formulation yields:
$W^{\prime}_{t+1}$ $=\alpha_{L}W_{L(t+1)}+\alpha_{S}W_{S(t+1)}$ (6)
$=\alpha_{L}\left(W_{L(t)}-\lambda_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}\right)+\alpha_{S}\left(W_{S(t)}-\lambda_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}\right)$ (7)
$={\alpha_{L}}W_{L(t)}+{\alpha_{S}}W_{S(t)}-{\lambda_{L}}{\alpha_{L}}{\frac{\partial{\mathcal{L}}}{\partial{W_{L(t)}}}}-{\lambda_{S}}{\alpha_{S}}{\frac{\partial{\mathcal{L}}}{\partial{W_{S(t)}}}}$ (8)
$=W^{\prime}_{t}-{\lambda_{L}}{\alpha_{L}}{\frac{\partial{\mathcal{L}}}{\partial{W_{L(t)}}}}-{\lambda_{S}}{\alpha_{S}}{\frac{\partial{\mathcal{L}}}{\partial{W_{S(t)}}}}$ (9)
From the equation 9, we observe that each branch can be optimized differently with different learning rates toward each kernel and derive with two distinctive scenarios as follow:
$W^{\prime}_{t+1}=\begin{cases}W^{\prime}_{t}-\lambda\left(\alpha_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}+\alpha_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}\right),&\text{if}\quad\lambda_{L}=\lambda_{S}\\ W^{\prime}_{t}-{\lambda_{L}}{\alpha_{L}}{\frac{\partial{\mathcal{L}}}{\partial{W_{L(t)}}}}-{\lambda_{S}}{\alpha_{S}}{\frac{\partial{\mathcal{L}}}{\partial{W_{S(t)}}}},&\text{if}\quad\lambda_{L}\neq\lambda_{S}\end{cases}$ (10)
By the chain rule, we further derive:
$\frac{\partial\mathcal{L}}{\partial W_{L(t)}}$ $=\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial Y_{\text{CSLA}}}{\partial W_{L(t)}}=\alpha_{L}\cdot\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial(X\ast W_{L})}{\partial W_{L(t)}},$ (11)
$\frac{\partial\mathcal{L}}{\partial W_{S(t)}}$ $=\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial Y_{\text{CSLA}}}{\partial W_{S(t)}}=\alpha_{S}\cdot\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial(X\ast W_{S})}{\partial W_{S(t)}}.$ (12)
To validate the above theoretical derivation, we perform ablation studies and found out that variable learning rate of each branch (i.e. $\lambda_{S}=0.0006,\lambda_{L}=0.0002$) demonstrates the best performance with stochastic gradient descent. Since $W_{S}$ has a smaller receptive field than $W_{L}$, and $W_{S}$ primarily contributes to the central region of the equivalent kernel $W^{\prime}$, we argue that:
  • Central region of $W^{\prime}$: receives gradient contributions from both $W_{L}$ and $W_{S}$, resulting in faster convergence and stronger local learning.
  • Peripheral region of $W^{\prime}$: receives gradients only from $W_{L}$, leading to slower convergence but maintaining global contextual awareness.
Both the coefficients $\alpha_{L}$ and $\alpha_{S}$ modulate spatially distinct regions (large kernel contributions dominate the periphery, small kernel contributions dominate the central region), the two–branch block demonstrates a learning-rate field of:
$\lambda_{\mathrm{eff}}(\Delta x)=\begin{cases}\alpha_{L}\,\lambda_{L},&\text{peripheral offsets},\\[4.0pt] \alpha_{L}\,\lambda_{L}+\alpha_{S}\,\lambda_{S},&\text{central offsets},\end{cases}$ (13)
where $\lambda_{eff}$ is the effective element-wise learning rate inherited from the two branch-specific updates.
3.2 Low-Rank Receptive Bias Modeling (LRBM)
As the above theory further validates the correlation between variable learning with the local-to-global gradient dynamics in ERF, we argue that such receptive bias can enhance the efficiency of learning large convolution kernels. We model the diffusion behavior of ERF with a reciprocal distance decay function $f_{d}$ and generate a prior mapping $P\in\mathbb{R}^{C\times 1\times K\times K\times K}$ for weight re-parameterization as follow:
$f_{d}(x,y,z,c)$ $=\sqrt{\left({x-c}\right)^{2}+\left({y-c}\right)^{2}+\left({z-c}\right)^{2}}$ (14)
$P$ $=\frac{\beta}{d(x_{k},y_{k},z_{k},c)+\beta}$
where $k$ and $c$ are the element and central index of the kernel weight, $\beta$ is a learnable parameter to control the weight distribution of the distance mapping and initialize as 0. However, such a fixed prior mapping lacks of flexibility to adapt the weighting importance dynamically across the fine-grained semantic variations in medical imaging. To address this, we propose to adapt learnable spatial bias by co-training a light-weight 2-layer generator network $f_{\theta}:\mathbb{R}^{C\times 1\times K\times K\times K}\rightarrow\mathbb{R}^{C\times 1\times K\times K\times K}$. We generate an adaptive mask $M$ for depthwise convolution kernels with low computation cost as follows:
$M=P+f_{\theta}(P)$ (15)
$f_{\theta}(P)=\text{Norm}_{2}(\text{DConv}_{2}\left(\sigma\left(\text{Norm}_{1}(\text{DConv}_{1}(P))\right)\right))$ (16)
where $\text{DConv}_{1}$ and $\text{DConv}_{2}$ are 3D depthwise convolutions with kernel size of 7 and padding of 3, both $\text{Norm}_{1}$ $\text{Norm}_{2}$ are the layer normalizations, and $\sigma$ is a non-linear sigmoid activation to ensure all scaling value between 0 and 1. Such learnable function aims to capture the dynamic weighting of each kernel elements across local to global, while preserving computational efficiency. The resulting modulation mask $M$ is then used to reparameterize the kernel weights:
$W_{\text{eff}}=W\odot M$ (17)
where $W$ is the original convolution kernel and $\odot$ denotes element-wise multiplication. Importantly, the mask is applied during training only and the learned generator can be removed during inference for efficiency.
3.3 Network Architecture
The overall network architecture to validate Rep3D builds upon the encoder–decoder structure of 3D UX-Net [18], which processes volumetric data through hierarchical resolution stages with skip connections to preserve fine-grained spatial features. Unlike prior transformer-based models or heavily modular CNNs, our design favors plain convolution blocks to minimize computational burden while preserving capacity for large-scale context modeling. Following the insights from prior work [7], we adopt a $21\times 21\times 21$ depthwise convolution (DWC-21) as the kernel backbone, which we empirically identify as the best trade-off between expressiveness and efficiency in 3D. Each encoder block consists of batch normalization, followed by the depthwise convolution and GeLU activation. The feature propagation from layer $\ell-1$ to $\ell$ and then to $\ell+1$ is defined as:
$\hat{z}_{\ell}=\text{GELU}\left(\text{DWC-21}(\text{BN}(z_{\ell-1}))\right),\quad\hat{z}_{\ell+1}=\text{GELU}\left(\text{DWC-21}(\text{BN}(\hat{z}_{\ell}))\right)$ (18)
where $z_{\ell-1}$ is the input from the previous layer, $\hat{z}\ell$ and $\hat{z}{\ell+1}$ are intermediate representations, BN denotes batch normalization, and DWC-21 represents depthwise convolution with a $21^{3}$ kernel. This architectural choice allows the network to efficiently encode both local and global context, while enabling seamless integration of our re-parameterized learning framework (as detailed in section 3.1 and 3.2). The simplicity of the block ensures compatibility with the spatial modulation mask described in the next section and avoids unnecessary overhead during both training and inference.
4 Experimental Setup
Datasets. We evaluate Rep3D on four publicly available volumetric segmentation datasets, covering a wide range of anatomical structures across different spatial scales—from large organs (e.g., liver, stomach) to smaller and more challenging targets (e.g., tumors, vessels): 1) AMOS22 (MICCAI 2022 Abdominal Multi-organ Segmentation Challenge) [15]: Comprises 200 multi-contrast abdominal CT scans with 15 organ-level anatomical labels and 33 MRI scans with 13 organ-level anatomical labels for comprehensive abdominal segmentation, 2) KiTS19 (MICCAI 2019 Kidney Tumor Segmentation Challenge) [12]: Includes 210 contrast-enhanced abdominal CT scans from the University of Minnesota Medical Center (2010–2018), with manual annotations for kidney, tumor, and cyst, 3) MSD Pancreas (Medical Segmentation Decathlon) [1]: Contains 282 abdominal contrast-enhanced CT scans annotated for both pancreas and pancreatic tumor segmentation, and 4) MSD Hepatic Vessel (Medical Segmentation Decathlon) [1]: Contains 303 abdominal CT scans annotated for hepatic vessel and associated tumor segmentation. Additional dataset details, including resolution normalization, voxel spacing, and pre-processing pipelines, are provided in the supplementary material.
Implementation Details. All experiments are conducted under a direct supervised learning setting. For the KiTS and MSD datasets, we employ a 5-fold cross-validation strategy using an 80%/10%/10% split for training, validation, and testing, respectively. For the AMOS dataset, we use a fixed single split with the same partitioning ratio. Details on training procedures and preprocessing protocols are provided in the supplementary material. Our proposed re-parameterization approach Rep3D, is benchmarked against both convolutional and transformer-based state-of-the-art (SOTA) methods for 3D medical image segmentation. For nnUNet [14], we evaluate performance across two training schedules to account for fairness, since Rep3D is trained with 60,000 iterations (approximately equivalent to 350 epochs). We report results using the Dice Similarity Coefficient (DSC) as the primary evaluation metric, quantifying spatial overlap between predicted segmentations and ground truth labels. Additionally, we conduct ablation studies to analyze: (1) the effect of kernel size in the spatial bias generator, and (2) the adaptability of Rep3D when integrated into other 3D large kernel architectures (e.g., 3D UX-Net).
Table 1: Comparison of SOTA approaches on the three different testing datasets. (*: $p<0.01$, with Paired Wilcoxon signed-rank test to all baseline networks)
KiTS MSD Methods #Params FLOPs Kidney Tumor Cyst Mean Pancreas Tumor Mean Hepatic Tumor Mean 3D U-Net [5] 4.81M 135.9G 0.918 0.657 0.361 0.645 0.711 0.584 0.648 0.569 0.609 0.589 SegResNet [26] 1.18M 15.6G 0.935 0.713 0.401 0.683 0.740 0.613 0.677 0.620 0.656 0.638 RAP-Net [20] 38.2M 101.2G 0.931 0.710 0.427 0.689 0.742 0.621 0.682 0.610 0.643 0.627 nn-UNet [14] 31.2M 743.3G 0.943 0.732 0.443 0.706 0.775 0.630 0.703 0.623 0.695 0.660 TransBTS [28] 31.6M 110.3G 0.932 0.691 0.384 0.669 0.749 0.610 0.679 0.589 0.636 0.613 UNETR [11] 92.8M 82.5G 0.921 0.669 0.354 0.648 0.735 0.598 0.667 0.567 0.612 0.590 nnFormer [32] 149.3M 213.0G 0.930 0.687 0.376 0.664 0.769 0.603 0.686 0.591 0.635 0.613 SwinUNETR [10] 62.2M 328.1G 0.939 0.702 0.400 0.680 0.785 0.632 0.708 0.622 0.647 0.635 3D UX-Net (k=7) [18] 53.0M 639.4G 0.942 0.724 0.425 0.697 0.737 0.614 0.676 0.625 0.678 0.652 UNesT-B [31] 87.2M 258.4G 0.943 0.746 0.451 0.710 0.778 0.601 0.690 0.611 0.645 0.640 Rep3D (Fixed Prior) 65.8M 757.4G 0.950 0.757 0.473 0.727 0.789 0.640 0.715 0.635 0.681 0.658 Rep3D 66.0M 757.6G 0.955 0.763 0.490 0.736* 0.793 0.653 0.723* 0.650 0.697 0.674*
Table 2: Evaluations on the AMOS testing split in different scenarios.(*: $p<0.01$, with Paired Wilcoxon signed-rank test to all baseline networks)
AMOS CT (Train From Scratch Scenario) Methods Spleen R. Kid L. Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg nn-UNet (350 Epochs) 0.951 0.919 0.930 0.845 0.797 0.975 0.863 0.941 0.898 0.813 0.730 0.677 0.772 0.797 0.815 0.850 nn-UNet (1000 Epochs) 0.967 0.958 0.945 0.890 0.818 0.979 0.914 0.953 0.920 0.824 0.799 0.743 0.823 0.900 0.867 0.887 TransBTS 0.930 0.921 0.909 0.798 0.722 0.966 0.801 0.900 0.820 0.702 0.641 0.550 0.684 0.730 0.679 0.783 UNETR 0.925 0.923 0.903 0.777 0.701 0.964 0.759 0.887 0.821 0.687 0.688 0.543 0.629 0.710 0.707 0.740 nnFormer 0.932 0.928 0.914 0.831 0.743 0.968 0.820 0.905 0.838 0.725 0.678 0.578 0.677 0.737 0.596 0.785 SwinUNETR 0.956 0.957 0.949 0.891 0.820 0.978 0.880 0.939 0.894 0.818 0.800 0.730 0.803 0.849 0.819 0.871 3D UX-Net (k=7) 0.966 0.959 0.951 0.903 0.833 0.980 0.910 0.950 0.913 0.830 0.805 0.756 0.846 0.897 0.863 0.890 3D UX-Net (k=21) 0.963 0.959 0.953 0.921 0.848 0.981 0.903 0.953 0.910 0.828 0.815 0.754 0.824 0.900 0.878 0.891 UNesT-B 0.966 0.961 0.956 0.903 0.840 0.980 0.914 0.947 0.912 0.838 0.803 0.758 0.846 0.895 0.854 0.891 RepOptimizer 0.968 0.964 0.953 0.903 0.857 0.981 0.915 0.950 0.915 0.826 0.802 0.756 0.813 0.906 0.867 0.892 Rep3D (Fixed Prior) 0.972 0.963 0.964 0.911 0.861 0.982 0.921 0.956 0.924 0.837 0.818 0.777 0.831 0.916 0.879 0.902 Rep3D (LRBM) 0.978 0.970 0.964 0.928 0.871 0.984 0.927 0.960 0.930 0.851 0.828 0.784 0.850 0.920 0.881 0.910* AMOS MRI (Train From Scratch Scenario) Methods Spleen R. Kid L. Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg nn-UNet (350 Epochs) 0.967 0.855 0.958 0.663 0.736 0.973 0.888 0.956 0.907 0.793 0.533 0.572 0.668 - - 0.805 nn-UNet (1000 Epochs) 0.973 0.940 0.965 0.681 0.810 0.980 0.893 0.967 0.917 0.834 0.667 0.689 0.701 - - 0.847 TransBTS 0.956 0.957 0.955 0.619 0.770 0.974 0.867 0.958 0.852 0.836 0.591 0.630 0.648 - - 0.816 UNETR 0.942 0.956 0.930 0.552 0.741 0.967 0.836 0.947 0.829 0.815 0.564 0.621 0.624 - - 0.794 nnFormer 0.949 0.952 0.950 0.601 0.758 0.972 0.859 0.960 0.843 0.832 0.569 0.618 0.637 - - 0.808 SwinUNETR 0.972 0.961 0.961 0.649 0.814 0.978 0.889 0.961 0.862 0.854 0.659 0.649 0.664 - - 0.836 3D UX-Net (k=7) 0.971 0.965 0.966 0.603 0.828 0.978 0.869 0.962 0.878 0.837 0.696 0.689 0.696 - - 0.841 3D UX-Net (k=21) 0.968 0.962 0.967 0.610 0.830 0.977 0.858 0.954 0.880 0.829 0.701 0.697 0.700 - - 0.840 UNesT-B 0.971 0.965 0.967 0.615 0.831 0.980 0.865 0.949 0.883 0.845 0.852 0.700 0.697 - - 0.854 RepOptimizer 0.970 0.967 0.971 0.635 0.823 0.978 0.875 0.963 0.882 0.850 0.689 0.691 0.711 - - 0.847 Rep3D (Fixed Prior) 0.972 0.965 0.970 0.644 0.838 0.980 0.883 0.965 0.893 0.861 0.714 0.701 0.725 - - 0.855 Rep3D (LRBM) 0.975 0.969 0.975 0.657 0.845 0.984 0.891 0.970 0.901 0.879 0.718 0.721 0.750 - - 0.864*
5 Results
5.1 Evaluation on Tissue & Tumor Segmentation
To assess the generalization and scalability of Rep3D across diverse anatomical structures and clinical targets, we evaluate performance on three representative volumetric segmentation tasks using the KiTS, MSD Pancreas, and MSD Hepatic Vessel datasets. As shown in Table 3, Rep3D achieves state-of-the-art performance across all settings, consistently outperforming both convolution- and transformer-based baselines. On the KiTS dataset, which includes kidney, tumor, and cyst segmentation, Rep3D achieves the highest average Dice score of 0.736, with strong individual scores of 0.955 (kidney), 0.763 (tumor), and 0.490 (cyst). Notably, Rep3D improves tumor segmentation performance by 2.28% Dice over UNesT-B and 5.39% Dice over 3D UX-Net, demonstrating its ability to adapt to complex local variations in pathological regions. On the MSD Pancreas task, which is particularly challenging due to the pancreas’s low contrast and irregular boundaries, Rep3D sets a new benchmark with an average Dice score of 0.723, outperforming SwinUNETR (0.708), nnUNet (0.703), and UNesT-B (0.690). Tumor segmentation also benefits from our re-parameterization design, improving by 3.32% Dice compared to 3D UX-Net and 2,03% Dice compared to the fixed-prior variant. On the MSD Hepatic Vessel dataset, Rep3D continues to lead with a mean Dice of 0.674, outperforming the previous best model (UNesT-B, 0.640) and demonstrating superior vessel and tumor localization. The results also highlight the effectiveness of Rep3D’s spatially adaptive learning dynamics, especially in sparse and small-structure segmentation where traditional large-kernel convolutions or global self-attention tend to underperform.
5.2 Evaluation on Multi-Organ Segmentation
Beyond the ability to segment anatomical structures across scales, we furtehr evaluate Rep3D on the AMOS benchmark under the "train-from-scratch" setting for both CT and MRI modalities. On AMOS-CT, Rep3D achieves the best performance across all 15 evaluated anatomical structures, surpassing strong baselines including SwinUNETR, UNesT, and 3D UX-Net. Notably, Rep3D outperforms UNesT-B by 2.13% and RepOptimizer by 2.02% of average Dice score, while operating with fewer parameters than UNesT. On AMOS-MRI, a more challenging modality due to the variable range of contrast intensity and anatomical ambiguity, Rep3D maintains its superior performance, achieving an average Dice of 0.864, again outperforming all competing approaches. Compared to the best-performing transformer baseline (UNesT-B, 0.854) and convolutional baseline (3D UX-Net (k=21), 0.840), Rep3D delivers consistent improvements across nearly all organ classes, particularly in difficult regions such as the pancreas, gallbladder, and adrenal glands. These gains underscore the effectiveness of our spatially adaptive re-parameterization strategy in enhancing convergence and feature expressivity without increasing model complexity.
Table 3: Ablation Studies on the AMOS testing split
Methods Spleen R. Kid L. Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg Kernel=1×1×$\times 1\times 1$ 1 0.972 0.968 0.965 0.926 0.863 0.984 0.917 0.956 0.922 0.851 0.816 0.779 0.863 0.912 0.894 0.905 Kernel=3×3×$\times 3\times 3$ 3 0.970 0.966 0.960 0.930 0.863 0.984 0.935 0.958 0.924 0.859 0.827 0.758 0.862 0.908 0.892 0.906 Kernel=5×5×$\times 5\times 5$ 5 0.974 0.967 0.964 0.925 0.833 0.984 0.924 0.956 0.910 0.850 0.829 0.786 0.843 0.921 0.884 0.903 Kernel=7×7×$\times 7\times 7$ 7 0.978 0.970 0.964 0.928 0.871 0.984 0.927 0.960 0.930 0.851 0.828 0.784 0.850 0.920 0.881 0.910 3D UX-Net (k=7) 0.966 0.959 0.951 0.903 0.833 0.980 0.910 0.950 0.913 0.830 0.805 0.756 0.846 0.897 0.863 0.890 3D UX-Net + LRBM 0.968 0.963 0.952 0.911 0.841 0.981 0.915 0.959 0.920 0.835 0.811 0.770 0.851 0.901 0.872 0.897
Figure 3: As kernel size increases, depthwise convolutions in 3D UX-Net exhibit increasingly diffuse ERFs, gradually expanding the gradient dynamics from local to broader spatial regions. Incorporating LRBM further enhances weighting toward global areas by modulating the spatial contribution of distant elements. In contrast, Rep3D produces a well-distributed ERF that preserves strong central activation while extending contextual influence across the full kernel.
5.3 Ablation Studies
Effect of Network Depth for LRBM. To investigate the impact of architectural depth in the spatial modulation generator, we conduct an ablation study by varying the number of layers in the generator network used to produce the element-wise modulation mask in Rep3D (in supplementart material). Specifically, we compare shallow configurations (1-layer depthwise convolution) with deeper variants (2-layer and 3-layer depthwise convolution stacks), while keeping the total parameter count approximately constant. Our results show that the 2-layer design provides the best trade-off between representation flexibility and training stability. While the 1-layer generator lacks sufficient capacity to capture nuanced spatial priors, resulting in under-modulated gradient flow. The 3-layer version demonstrates a slight decrease of performance (from 0.910 to 0.899 Dice) and instability during training. This suggests that a lightweight, moderately deep generator is optimal for learning spatially adaptive convergence patterns without incurring additional complexity or over-parameterization.
Effect of Kernel Size in Spatial Bias Modeling. To further understand how kernel size affects segmentation performance across different anatomical structures, we analyzed organ-wise performance under varying kernel configurations: $1\times 1\times 1$, $3\times 3\times 3$, $5\times 5\times 5$, and $7\times 7\times 7$ used in Rep3D. All configurations share the same training protocol and re-parameterization setup, isolating the effect of kernel size alone. As shown in Table 4, the impact of kernel size varies across organs. While the $7\times 7\times 7$ kernel achieves the highest overall mean Dice score (0.910), smaller or boundary-sensitive organs (e.g., bladder, adrenal glands) benefit from small- or mid-size kernels such as $1\times 1\times 1$ or $5\times 5\times 5$. In contrast, large organs with strong spatial continuity (e.g., liver, spleen, aorta) show clear improvements with larger receptive fields. These results suggest that optimal kernel size is organ-dependent, influenced by factors such as spatial extent, anatomical context, and structural complexity. The superior performance of the $7\times 7\times 7$ variant reflects its ability to balance local detail and global context.
Effect of LRBM towards Other Network Architectures. To isolate the contribution of our proposed Low-Rank Bias Modeling (LRBM) module, we integrate it into a standard 3D UX-Net architecture (with fixed $7\times 7\times 7$ kernels) and compare its performance to the original baseline. As reported in Table 4, incorporating LRBM improves the average Dice score from 0.890 to 0.897, with consistent gains across multiple organs including pancreas, bladder, and adrenal glands. While the improvement may appear modest in aggregate, it is particularly noteworthy in anatomically challenging regions where gradient convergence is often unstable. For example, performance on the left adrenal gland increases from 0.756 to 0.770, and the duodenum improves from 0.846 to 0.851, suggesting that the learnable spatial bias improves optimization dynamics in fine-scale structures. These results confirm that our LRBM module offers a generalizable and plug-and-play mechanism for enhancing 3D segmentation backbones, even outside the full Rep3D framework.
6 Discussions & Limitations
In this work, we introduced Rep3D, a re-parameterization framework that explicitly models spatial convergence dynamics in large kernel 3D convolutions. By linking effective receptive field (ERF) behavior with first-order optimization theory, we demonstrated that large convolution kernels naturally exhibit non-uniform learning dynamics, where central elements converge faster than peripheral ones. To address this, Rep3D integrates a learnable spatial prior via low-rank modulation, allowing the optimizer to differentially emphasize kernel regions with the distinctive characteristics of ERF during training. Our experiments across five diverse 3D segmentation benchmarks, confirm that Rep3D consistently improves performance over both transformer-based and convolution-based SOTA approaches, while maintaining a plain and efficient encoder design. The success of Rep3D reinforces several broader insights. First, spatially adaptive optimization is a promising direction for bridging inductive biases in CNNs with the dynamic learning capacity of attention-based models. Second, incorporating explicit ERF modeling into kernel design enables more efficient parameter usage, particularly in data-limited medical imaging scenarios. Moreover, our framework enhance network interpretability: the modulation masks can be visualized and aligned with ERF patterns, offering insights into how spatial understanding guides the learning of convolution kernels.
While Rep3D demonstrates strong empirical performance across diverse 3D medical segmentation tasks, several limitations remain. First, although our learnable modulation mechanism introduces minimal architectural overhead, the training cost associated with large 3D kernels (e.g., $21\times 21\times 21$) remains nontrivial, particularly in memory-constrained GPU environments. Unlike 2D convolution kernels (i.e. MegEngine packages for 2D depthwise kernels), limited packages and approaches has been proposed to optimize the large kernel mechanism in 3D. This limits the batch size and input resolution during training, which can affect convergence and generalization. Future work could explore progressive training strategies, multi-resolution optimization, or low-resolution proxy supervision to alleviate this constraint while maintaining segmentation fidelity. Second, while our distance decay prior effectively guides spatial re-parameterization, its performance is inherently tied to the input volume resolution. In our experiments, we downsample 3D volumes to specific resolution (e.g., $1.5\times 1.5\times 2.0$ mm) to balance computation and efficency. However, we observe saturation effects when training at higher resolutions, where further improvements in image quality do not yield proportional gains in segmentation accuracy. This may be due to the spatial prior losing precision at finer scales. Adapting fine-grained spatial learnable prior could be another potential direction for future work.
7 Conclusion
In this paper, we introduced Rep3D, a receptive-biased re-parameterization framework for large kernel 3D convolutions. By modeling effective receptive field (ERF) behavior as a learnable spatial prior, Rep3D enables adaptive element-wise learning dynamics during training, bridging the gap between convolutional inductive bias and optimization-aware design. Implemented via a lightweight modulation network, our approach avoids complex multi-branch architectures while improving training efficiency and segmentation accuracy. Extensive experiments across five volumetric medical imaging benchmarks demonstrate consistent improvements over SOTA transformer and CNN approaches, establishing Rep3D as a scalable and effective solution for 3D medical image analysis.
Appendix A Data Preprocessing & Training Details
Table 4: Hyperparameters for direction training scenario on four public datasets
Hyperparameters Direct Training Encoder Stage 4 Layer-wise Channel $48,96,192,384$ Hidden Dimensions $768$ Patch Size $96\times 96\times 96$ No. of Sub-volumes Cropped 2 Training Steps 60000 Batch Size 2 AdamW $\epsilon$ $81e-$ 8 AdamW $\beta$ ($0.9,0.999$) Peak Learning Rate $41e-$ 4 Learning Rate Scheduler ReduceLROnPlateau Factor & Patience 0.9, 10 Dropout X Weight Decay 0.08 Data Augmentation Intensity Shift, Rotation, Scaling Cropped Foreground Intensity Offset 0.1 Rotation Degree $-30^{\circ}$ to $+30^{\circ}$ Scaling Factor x: 0.1, y: 0.1, z: 0.1
We apply hierarchical steps for data preprocessing: 1) intensity clipping is applied to further enhance the contrast of soft tissue (AMOS CT, KiTS, MSD Pancreas:{min:-175, max:250}; MSD Hepatic Vessel:{min:0, max:230}); AMOS MRI:{min:0, max:1000}. 2) Intensity normalization is performed after clipping for each volume and use min-max normalization: $(X-X_{1})/(X_{99}-X_{1})$ to normalize the intensity value between 0 and 1, where $X_{p}$ denote as the $p_{th}$ percentile of intensity in $X$. We then perform downsampling to certain voxel spacing (i.e. AMOS CT, MSD hepatic vessels, MSD Pancreas and KiTS: $1.5\times 1.5\times 2.0$, AMOS MRI: $1.0\times 1.0\times 1.0$) randomly crop sub-volumes with size $96\times 96\times 96$ at the foreground and perform data augmentations, including rotations, intensity shifting, and scaling (scaling factor: 0.1). All training processes with Rep3D are optimized with either Stochastic Gradient Descent (SGD) or AdamW optimizer. We trained all models for 60000 steps using a learning rate of $0.0001$ on an NVIDIA A100 GPU across all datasets. One epoch takes approximately about 9 minute for KiTS, 5 minutes for MSD Pancreas, 12 minutes for MSD hepatic vessels, 7 minutes for AMOS CT and 1 minute for AMOS MRI, respectively. We further summarize all the training parameters with Table 1 in Supplementary Material.
Appendix B Datasets Details
Table 5: Complete overview of Four public datasets
Challenge AMOS CT AMOS MR MSD Pancreas MSD Hepatic Vessels KiTS Imaging Modality Multi-Contrast CT Multi-Contrast MRI Venous CT Arterial CT Anatomical Region Abdomen Pancreas Liver Kidney Sample Size 361 33 282 303 300 Anatomical Label Spleen, Left & Right Kidney, Gall Bladder, Pancreas, Tumor Hepatic Vessels, Tumor Kidney, Tumor Esophagus, Liver, Stomach, Aorta, Inferior Vena Cava (IVC) Pancreas, Left & Right Adrenal Gland (AG), Duodenum Bladder (CT only), Prostate/Uterus (CT only) Data Splits 1-Fold (Internal) 5-Fold Cross-Validation Train: 160 / Validation: 20 / Test: 20 Train: 22 / Validation: 4/ Test: 7 Train: 225 / Validation: 27 / Testing: 30 Training: 242, Validation: 30 / Testing: 31 Training: 240, Validation: 30 / Testing: 30 5-Fold Ensembling N/A N/A X X
Appendix C Network Architecture
We adopt a 3D encoder-decoder architecture from both 3D UX-Net [18] and SwinUNETR [10] as the backbone of Rep3D. Instead of using encoder block with feed forward layer, we simply using a plain convolutional design with depthwise separable convolutions in parallel with LRBM. The encoder consists of 4 hierarchical stages with increasing feature dimensions and depthwise convolutions of large kernel size ($21\times 21\times 21$), followed by a symmetric decoder for volumetric segmentation. The encoder includes:
  • An initial input projection block with a $7\times 7\times 7$ convolution (stride 2, padding 3) followed by a residual block with two $3\times 3\times 3$ convolutions and GELU activations.
  • Stage 1: 2 Rep3D blocks with 48 channels followed by a strided $2\times 2\times 2$ convolution for downsampling.
  • Stage 2: 2 Rep3D blocks with 96 channels, followed by a strided $2\times 2\times 2$ convolution for downsampling.
  • Stage 3: 2 Rep3D blocks with 192 channels, followed by a strided $2\times 2\times 2$ convolution for downsampling.
  • Stage 4: 2 Rep3D blocks with 384 channels, followed by a strided $2\times 2\times 2$ convolution for downsampling.
Each stage modulates large kernel weights using a learnable re-parameterization mask computed via a lightweight 2-layer generator network within each Rep3D block. For each Rep3D block, it includes:
  • A single depthwise 3D convolution with a large kernel size of $21\times 21\times 21$ and padding size of $10$, followed by a layer normalization and a GELU activation.
  • A 2-stage lightweight generator network including:
    • First layer: a depthwise $7\times 7\times 7$ convolution followed by layer normalization and a sigmoid activation.
    • Second layer: another depthwise $7\times 7\times 7$ convolution followed by layer normalization.
The decoder mirrors the encoder and consists of:
  • 4 upsampling modules (UnetrUpBlock from MONAI), each with a transpose convolution (stride 2), skip connection, and a residual block with two $3\times 3\times 3$ convolutions and GELU activations.
  • 1 output projection block (UnetOutBlock from MONAI) consisting of a $1\times 1\times 1$ convolution to map to the number of target classes.
Appendix D Validation Experiments on Variable Branch Learning rate
Table 6: Quantitative Evaluation on Variable Learning Rates in Parallel Branches
Optimizer  Main Branch Para. Branch  Train Steps Main LR Para. LR Mean Dice SGD $21\times 21\times 21$$\times$ 60000 0.0005 $\times$ 0.849 SGD $21\times 21\times 21$$\times$ 60000 0.0004 $\times$ 0.852 SGD $21\times 21\times 21$$\times$ 60000 0.0003 $\times$ 0.856 SGD $21\times 21\times 21$$\times$ 60000 0.0002 $\times$ 0.859 SGD $21\times 21\times 21$$\times$ 60000 0.0001 $\times$ 0.854 AdamW $21\times 21\times 21$$\times$ 60000 0.0005 $\times$ 0.855 AdamW $21\times 21\times 21$$\times$ 60000 0.0004 $\times$ 0.859 AdamW $21\times 21\times 21$$\times$ 60000 0.0003 $\times$ 0.861 AdamW $21\times 21\times 21$$\times$ 60000 0.0002 $\times$ 0.862 AdamW $21\times 21\times 21$$\times$ 60000 0.0001 $\times$ 0.860 SGD $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0006 0.872 SGD $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0005 0.869 SGD $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0004 0.867 SGD $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0003 0.870 SGD $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0001 0.865 AdamW $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0006 0.887 AdamW $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0005 0.886 AdamW $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0004 0.887 AdamW $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0003 0.889 AdamW $21\times 21\times 21$ $3\times 3\times 3$ 60000 0.0002 0.0001 0.886
To empirically validate the theoretical insight of the spatially varying convergence dynamics in parallel-branched re-parameterization, we initially perform experiments using the CSLA block with Rep3D network architecture, composing of a main large kernel branch ($21\times 21\times 21$) and a parallel small kernel branch ($3\times 3\times 3$), with separate learning rates applied to each. As shown in Table 3, the single-branch design (no parallel branch) performance improved moderately with lower learning rates with both SGD and AdamW. The Dice score peaks at 0.859 with a learning rate of 0.0002 using SGD, and AdamW achieves its best performance of 0.862 at 0.0002 as well. However, with the addition of a small kernel parallel branch and using a higher learning rate for the small kernel (e.g., $\lambda_{\text{S}}>\lambda_{\text{L}}$), we observed consistent improvements across all configurations. Specifically, the best result with SGD reached 0.872 when using $\lambda_{L}=0.0002$ and $\lambda_{S}=0.0006$. Similarly, AdamW attained a maximum Dice score of 0.889 with $\lambda_{L}=0.0002$ and $\lambda_{S}=0.0003$. These results validate our hypothesis that assigning higher learning rates to the small kernel branch accelerates convergence of central kernel regions, while maintaining stability in peripheral regions with a lower learning rate for the large kernel. Moreover, such results further confirm that spatially varying convergence behavior can be approximated through differentiated learning rates, supporting the design principle behind our learnable re-parameterization in Rep3D.
Appendix E Ablation Study on Network Depth for LRBM
Table 7: Ablation Study on Network Depth for LRBM with the AMOS testing split
Number of Layers Spleen R. Kid L. Kid Gall. Eso. Liver Stom. Aorta IVC Panc. RAG LAG Duo. Blad. Pros. Avg 1 Layer 0.974 0.965 0.964 0.925 0.859 0.982 0.926 0.956 0.920 0.842 0.824 0.781 0.842 0.915 0.879 0.904 2 Layers 0.978 0.970 0.964 0.928 0.871 0.984 0.927 0.960 0.930 0.851 0.828 0.784 0.850 0.920 0.881 0.910 3 Layers 0.971 0.964 0.965 0.924 0.841 0.983 0.920 0.952 0.910 0.839 0.819 0.779 0.837 0.910 0.870 0.899